# Graph-Based Fault Diagnosis for Rotating Machinery: Adaptive Segmentation and Structural Feature Integration

### v1

@@ 3.2 Features Extraction
$\gamma_{X_{i}}=\frac{\mathbb{E}\left[\left(\frac{1}{2}\sum_{k=1}^{ch}(X_{i,j,k}-\mu_{X_{i}})\right)^{3}\right]}{\sigma_{X_{i}}^{3}}$
$\kappa_{X_{i}}=\frac{\mathbb{E}\left[\left(\frac{1}{2}\sum_{k=1}^{ch}(X_{i,j,k}-\mu_{X_{i}})\right)^{4}\right]}{\sigma_{X_{i}}^{4}}.$
@@
$\text{RMS}_{X_{i}}=\sqrt{\frac{1}{2N}\sum_{k=1}^{ch}\sum_{j=1}^{N}X_{i,j,k}^{2}}$
$\text{Peak}_{X_{i}}=\max_{k,j}|X_{i,j,k}|$
@@
$\text{TKEO}_{X_{i},k}=\frac{1}{N-2}\sum_{j=1}^{N-2}(X_{i,j+1,k}^{2}-X_{i,j,k}X_{i,j+2,k})$
$\text{TKEO}_{X_{i}}=\frac{1}{2}\sum_{k=1}^{ch}\text{TKEO}_{X_{i},k}$
@@
$P_{X_{i}X_{i}}(f)=\frac{1}{K_{w}}\sum_{m=1}^{K_{w}}|\mathcal{F}\{\mathbf{X}_{i,m}\}|^{2}$
@@
$A_{X_{i},k}(t)$ =|$=\left|\mathcal{H}\left\{\mathbf{X}_{i,:,k}(t)\right\}\right|$
[…]
@@ 3.3 K-Nearest Neighbor Graph Construction
$\mathcal{E}_{G}=\{(i,j)\mid j\in\text{Top-}k(\text{NN}(\mathbf{F}_{i})),\,d(\mathbf{F}_{i},\mathbf{F}_{j})<\tau\},$
@@
$w_{ij}=\frac{1}{1+\|\mathbf{F}_{i}-\mathbf{F}_{j}\|_{2}}.$
To improve computational efficiency, we leveraged the MiniBatch K-Means clustering results from the previous section, reducing the number of pairwise distance comparisons by prioritizing connections within and between clusters. Large clusters were decomposed into modular subgraphs.
$\mathcal{G}=\{\mathcal{G}_{1},\mathcal{G}_{2},\dots,\mathcal{G}_{m_{G}}\},\quad|\mathcal{G}_{l}|\leq N_{\text{max}},$
@@ 3.4 Graph-Theoretic Features
Each node $v\in\mathcal{V}_{l}$ corresponds to a segment and carries a feature vector $\mathbf{F}_{v}\in\mathbb{R}^{d}$ that encodes local properties such as time-frequency characteristics, along with a multiclass fault label $\ell_{v}\in\{0,1,\ldots,K_{F}-1\}$, where $\ell_{v}=0$ represents normal operation, and $11,\ldots,K_{F}-$ 1 corresponds to distinct fault types.
@@
$L_{\mathcal{G}_{l}}=\frac{1}{|\mathcal{V}_{l}|(|\mathcal{V}_{l}|-1)}\sum_{u\neq v\in\mathcal{V}_{l}}d_{\text{path}}(u,v),$
@@
$d_{\text{path}}(u,v)=\min_{\text{path }p:u\to v}\sum_{(s,t)\in p}w_{st},$
@@
$L_{\text{avg}}=\frac{1}{|\mathcal{C}_{\text{conn}}|}\sum_{\mathcal{G}_{l}\in\mathcal{C}_{\text{conn}}}L_{\mathcal{G}_{l}},$
@@
$d_{\text{avg}}=\frac{1}{m_{G}}\sum_{\mathcal{G}_{l}}\frac{2|\mathcal{E}_{\mathcal{G}_{l}}|}{|\mathcal{V}_{l}|},$
@@
$L_{\text{exp-sw}}=\frac{\log N_{\text{total}}}{\log d_{\text{avg}}},\quad N_{\text{total}}=\sum_{l}|\mathcal{V}_{l}|,$
@@
$Q_{\mathcal{G}_{l}}=\frac{1}{2|\mathcal{E}_{\mathcal{G}_{l}}|}\sum_{u,v\in\mathcal{V}_{l}}\left[A_{\mathcal{G}_{l}}(u,v)-\frac{k_{u}k_{v}}{2|\mathcal{E}_{\mathcal{G}_{l}}|}\right]\delta(\ell_{u},\ell_{v}),$
@@
$Q_{\text{mod}}=\frac{1}{|\mathcal{E}_{\text{nonempty}}|}\sum_{\mathcal{G}_{l}\in\mathcal{E}_{\text{nonempty}}}Q_{\mathcal{G}_{l}},$
@@
$\mathbf{L}_{\mathcal{G}_{l}}(u,v)=\begin{cases}-w_{uv}&\text{if }(u,v)\in\mathcal{E}_{\mathcal{G}_{l}},\\ \sum_{z\in\mathcal{V}_{l}}w_{uz}&\text{if }u=v,\\ 0&\text{otherwise}.\end{cases}$
@@
$\Delta_{\text{spec}}=\frac{1}{|\mathcal{C}_{\text{conn}}|}\sum_{\mathcal{G}_{l}\in\mathcal{C}_{\text{conn}}}\Delta_{\mathcal{G}_{l}}.$
@@ 3.5 Multiclass Fault Detection Model
Finally, we built a feature set for multiclass fault detection by combining local and global information. For each node $v$, we define the feature vector as
@@
where $\mathbf{F}_{v}$ captures node-specific features (e.g., time-frequency characteristics), and $L_{\text{avg}}$, $Q_{\text{mod}}$, and $\Delta_{\text{spec}}$ represent system-level structural properties derived from graph-theoretical analysis. The corresponding label is $\ell_{v}^{\prime}=\ell_{v}$, where $\ell_{v}\in\{0,1,\ldots,K-1\}$ was defined previously. These are aggregated into a feature matrix $\mathbf{Z}\in\mathbb{R}^{N_{\text{total}}\times(d+3)}$ and label vector $\mathbf{y}_{\text{label}}\in\{0,1,\ldots,K-1\}^{N_{\text{total}}}$, where $N_{\text{total}}$ is the total number of nodes across all the subgraphs.
Logistic regression models the probability of fault class $\ell_{v}=k$ given feature vector $\mathbf{z}_{v}$:
$P(\ell_{v}=k\mid\mathbf{z}_{v};\mathbf{W},\mathbf{b})=\frac{\exp(\mathbf{w}_{k}^{\top}\mathbf{z}_{v}+b_{k})}{\sum_{j=0}^{K-1}\exp(\mathbf{w}_{j}^{\top}\mathbf{z}_{v}+b_{j})}$ (5)
where $\mathbf{w}_{k}\in\mathbb{R}^{d+3}$ and $b_{k}\in\mathbb{R}$ are the weight vector and bias for class $k$, $\mathbf{W}=[\mathbf{w}_{0},\ldots,\mathbf{w}_{K-1}]$ is the weight matrix, and $\mathbf{b}=[b_{0},\ldots,b_{K-1}]$ is the bias vector. The model is trained by minimizing the cross-entropy loss:
$L(\mathbf{W},\mathbf{b})=-\frac{1}{N}\sum_{i=1}^{N}\sum_{k=0}^{K-1}\mathbb{1}\{\ell_{i}=k\}\log P(\ell_{i}=k\mid\mathbf{z}_{i}),$ (6)
using gradient descent. The predicted class is:
^$\hat{\ell}_{v}=\arg\max_{k}P(\ell_{v}=k\mid\mathbf{z}_{v}).$
@@
Transfer learning adapts a model trained on a source domain $\mathcal{D}_{S}=\{(\mathbf{z}_{i}^{S},\ell_{i}^{S})\}_{i=1}^{N_{S}}$ to a target domain $\mathcal{D}_{T}=\{(\mathbf{z}_{i}^{T},\ell_{i}^{T})\}_{i=1}^{N_{T}}$. Both domains share the same feature space (e.g., $\mathbf{z}_{v}$ with FFT, modularity) and label space ($\{0,\ldots,K-1\}$) but differ in distributions due to domain shifts (e.g., load-induced FFT frequency changes from 50–100 Hz to 55–95 Hz (Zhang et al., 2017; Chennana et al., 2025)).
The process comprises:
  1. 1.
    Source Training: Train logistic regression on $\mathcal{D}_{S}$ to obtain optimal parameters $\mathbf{W}_{S},\mathbf{b}_{S}$ by minimizing $L(\mathbf{W},\mathbf{b})$. This captures general fault patterns.
  2. 2.
    Weight Transfer: Initialize the target model with $\mathbf{W}_{S},\mathbf{b}_{S}$, leveraging learned feature-fault relationships.
  3. 3.
    Fine-Tuning: Fine-tune on $\mathcal{D}_{T}$ by minimizing the target loss $L(\mathbf{W},\mathbf{b})$ via gradient descent, starting from $\mathbf{W}_{S},\mathbf{b}_{S}$. This adapts weights to target-specific patterns.
The loss gradient for a target sample $(\mathbf{z}_{i}^{T},\ell_{i}^{T})$ is:
$\nabla_{\mathbf{w}_{k}}L=\frac{1}{N_{T}}\sum_{i=1}^{N_{T}}\left[P(\ell_{i}=k\mid\mathbf{z}_{i}^{T})-\mathbb{1}\{\ell_{i}=k\}\right]\mathbf{z}_{i}^{T},$
$\nabla_{b_{k}}L=\frac{1}{N_{T}}\sum_{i=1}^{N_{T}}\left[P(\ell_{i}=k\mid\mathbf{z}_{i}^{T})-\mathbb{1}\{\ell_{i}=k\}\right].$
@@ 3.6 Algorithmic Overview and Complexity Analysis
The proposed method for multiclass fault detection comprises five key processes: feature extraction from signal segments, segmentation and clustering of feature vectors, KNN graph construction, graph-theoretic feature extraction, and multiclass fault detection.
@@
Feature extraction begins by segmenting the $ch$-channel vibration signal $\mathbf{S}\in\mathbb{R}^{N\times ch}$ into $n=\lfloor(N-\omega^{*})/\sigma^{*}\rfloor$ segments using the optimal window size $\omega^{*}$ and step size $\sigma^{*}$. For each segment, time- and frequency-domain features (e.g., mean, RMS, FFT, Hilbert transform (Welch, 1967; Hilbert, 1930)) are computed per channel, yielding a feature vector of dimension $d=ch\cdot d_{c}$, where $d_{c}$ is the number of features per channel (e.g., 20 features per channel for $ch=2$ yields $d=40$ (Chennana et al., 2025)). The dominant operation is FFT, with complexity $O(\omega^{*}\log\omega^{*})$ per segment per channel, resulting in a total complexity of $O(chn\omega^{*}\log\omega^{*})$.
Next, segmentation and clustering employ MiniBatch K-Means to group $F$ feature vectors of dimension $d$ into $K_{c}$ clusters, with a complexity of $O(FdK_{c})=O(Fchd_{c}K_{c})$, driven by distance computations over $T$ iterations and batch size $B$. KNN graph construction builds a graph per cluster, connecting each node to its $k_{\text{NN}}$ nearest neighbors within a threshold $\tau$, and decomposes large clusters into subgraphs of size at most $N_{\text{max}}$. With approximately $F/K_{c}$ nodes per cluster, distance computations and neighbor selection require $O((F/K_{c})(d+\log(F/K_{c})))=O((F/K_{c})(chd_{c}+\log(F/K_{c})))$ per node, leading to a total complexity of $O(F(F/K_{c})(chd_{c}+\log(F/K_{c})))$, which dominates the pipeline.
Graph-theoretic feature extraction computes the average shortest path length $L_{\text{avg}}$, modularity $Q_{\text{mod}}$, and spectral gap $\Delta_{\text{spec}}$ (Newman, 2006; Cheeger, 1970) for each of the $m_{G}\approx K_{c}$ subgraphs, where each subgraph has $|V_{l}|\approx N_{\text{total}}/m_{G}$ nodes ($N_{\text{total}}\approx n$). The dominant operation, Dijkstra’s algorithm for shortest paths, takes $O(|V_{l}|^{2}\log|V_{l}|)$ per subgraph, yielding a total complexity of $O((N_{\text{total}}^{2}/m_{G})\log(N_{\text{total}}/m_{G}))$.
Finally, multiclass fault detection constructs a feature matrix $\mathbf{Z}\in\mathbb{R}^{N_{\text{total}}\times(d+3)}$ $=$ $\mathbb{R}^{N_{\text{total}}\times(chd_{c}+3)}$, trains a softmax classifier over $T$ epochs, and predicts labels with a complexity of $O(TN_{\text{total}}K(chd_{c}+3))$, where $K$ is the number of fault classes. Assuming $\omega^{*}$, $d_{c}$, $K_{c}$, $k_{\text{NN}}$, $K$, and $T$ are constants, the KNN graph construction dominates with complexity $O((F^{2}/K_{c})(chd_{c}+\log(F/K_{c})))$, simplifying to $O(F^{2}/K_{c})$ for large $F$. The overall complexity remains $O(F^{2}/K_{c})$, which can be reduced by increasing $K_{c}$, using approximate nearest neighbor search, or parallelizing computations.

### v2

@@ 2.2 Features Extraction
$\gamma_{X_{i}}=\frac{\mathbb{E}\left[\left(\frac{1}{2}\sum_{k=1}^{ch}(X_{i,j,k}-\mu_{X_{i}})\right)^{3}\right]}{\sigma_{X_{i}}^{3}}$ (15)
$\kappa_{X_{i}}=\frac{\mathbb{E}\left[\left(\frac{1}{2}\sum_{k=1}^{ch}(X_{i,j,k}-\mu_{X_{i}})\right)^{4}\right]}{\sigma_{X_{i}}^{4}}.$ (16)
@@
$\text{RMS}_{X_{i}}=\sqrt{\frac{1}{2N}\sum_{k=1}^{ch}\sum_{j=1}^{N}X_{i,j,k}^{2}}$ (17)
$\text{Peak}_{X_{i}}=\max_{k,j}|X_{i,j,k}|$ (18)
@@
$\text{TKEO}_{X_{i},k}=\frac{1}{N-2}\sum_{j=1}^{N-2}(X_{i,j+1,k}^{2}-X_{i,j,k}X_{i,j+2,k})$ (19)
$\text{TKEO}_{X_{i}}=\frac{1}{2}\sum_{k=1}^{ch}\text{TKEO}_{X_{i},k}$ (20)
@@
$P_{X_{i}X_{i}}(f)=\frac{1}{K_{w}}\sum_{m=1}^{K_{w}}|\mathcal{F}\{\mathbf{X}_{i,m}\}|^{2}$ (21)
@@
$A_{X_{i},k}(t)$ $=\left|\mathcal{H}\left\{\mathbf{X}_{i,:,k}(t)\right\}\right|$ (22)
$=\sqrt{\mathbf{X}_{i,:,k}^{2}(t)+\left(\frac{1}{\pi}\,\text{p.v.}\int_{-\infty}^{\infty}\frac{\mathbf{X}_{i,:,k}(\tau)}{t-\tau}\,d\tau\right)^{2}}$
@@ 2.3 K-Nearest Neighbor Graph Construction
$\mathcal{E}_{G}=\{(i,j)\mid j\in\text{Top-}k(\text{NN}(\mathbf{F}_{i})),\,d(\mathbf{F}_{i},\mathbf{F}_{j})<\tau\}$ (24)
@@
$w_{ij}=\frac{1}{1+\|\mathbf{F}_{i}-\mathbf{F}_{j}\|_{2}}$ (25)
To improve computational efficiency, we leveraged the MiniBatch K-Means clustering results from the previous section, reducing the number of pairwise distance comparisons by prioritizing connections within and between clusters. Large clusters were decomposed into modular subgraphs
$\mathcal{G}=\{\mathcal{G}_{1},\mathcal{G}_{2},\dots,\mathcal{G}_{m_{G}}\},\quad|\mathcal{G}_{l}|\leq N_{\text{max}}$ (26)
@@ 2.4 Graph-Theoretic Features
Each node $v\in\mathcal{V}_{l}$ corresponds to a segment and carries a feature vector $\mathbf{F}_{v}\in\mathbb{R}^{d}$ that encodes local properties such as time-frequency characteristics, along with a multiclass fault label $\ell_{v}\in\{0,1,\ldots,K_{F}-1\}$, where $\ell_{v}=0$ represents normal operation, and $1,\ldots,K_{F}-1$ corresponds to distinct fault types.
@@
$L_{\mathcal{G}_{l}}=\frac{1}{|\mathcal{V}_{l}|(|\mathcal{V}_{l}|-1)}\sum_{u\neq v\in\mathcal{V}_{l}}d_{\text{path}}(u,v)$ (27)
@@
$d_{\text{path}}(u,v)=\min_{\text{path }p:u\to v}\sum_{(s,t)\in p}w_{st}$ (28)
@@
$L_{\text{avg}}=\frac{1}{|\mathcal{C}_{\text{conn}}|}\sum_{\mathcal{G}_{l}\in\mathcal{C}_{\text{conn}}}L_{\mathcal{G}_{l}}$ (29)
@@
$d_{\text{avg}}=\frac{1}{m_{G}}\sum_{\mathcal{G}_{l}}\frac{2|\mathcal{E}_{\mathcal{G}_{l}}|}{|\mathcal{V}_{l}|}$ (30)
@@
$L_{\text{exp-sw}}=\frac{\log N_{\text{total}}}{\log d_{\text{avg}}},\quad N_{\text{total}}=\sum_{l}|\mathcal{V}_{l}|$ (31)
@@
$Q_{\mathcal{G}_{l}}=\frac{1}{2|\mathcal{E}_{\mathcal{G}_{l}}|}\sum_{u,v\in\mathcal{V}_{l}}\left[A_{\mathcal{G}_{l}}(u,v)-\frac{k_{u}k_{v}}{2|\mathcal{E}_{\mathcal{G}_{l}}|}\right]\delta(\ell_{u},\ell_{v})$ (32)
@@
$Q_{\text{mod}}=\frac{1}{|\mathcal{E}_{\text{nonempty}}|}\sum_{\mathcal{G}_{l}\in\mathcal{E}_{\text{nonempty}}}Q_{\mathcal{G}_{l}}$ (33)
@@
$\mathbf{L}_{\mathcal{G}_{l}}(u,v)=\begin{cases}-w_{uv}&\text{if }(u,v)\in\mathcal{E}_{\mathcal{G}_{l}},\\ \sum_{z\in\mathcal{V}_{l}}w_{uz}&\text{if }u=v,\\ 0&\text{otherwise}.\end{cases}$ (34)
@@
$\Delta_{\text{spec}}=\frac{1}{|\mathcal{C}_{\text{conn}}|}\sum_{\mathcal{G}_{l}\in\mathcal{C}_{\text{conn}}}\Delta_{\mathcal{G}_{l}}$ (35)
@@ 2.5 Multiclass Fault Detection Model
To enable multiclass fault detection, we construct a comprehensive feature set that integrates both local and global information. Specifically, for each node $v$, we define a feature vector:
@@
where $\mathbf{F}_{v}$ captures node-level attributes, such as time-frequency characteristics, whereas $L_{\text{avg}}$, $Q_{\text{mod}}$, and $\Delta_{\text{spec}}$ are graph-theoretic metrics that encode system-level structural properties. The label for each node remains $\ell_{v}^{\prime}=\ell_{v}$, with $\ell_{v}\in\{0,1,\ldots,K-1\}$ indicating one of $K$ fault categories. These features and labels are assembled into a matrix $\mathbf{Z}\in\mathbb{R}^{N_{\text{total}}\times(d+3)}$ and a label vector $\mathbf{y}_{\text{label}}\in\{0,1,\ldots,K-1\}^{N_{\text{total}}}$, where $N_{\text{total}}$ denotes the total number of nodes in all subgraphs.
We then formulate a multiclass classification problem, where a model $f(\mathbf{z}_{v};\theta)$, parameterized by $\theta$, maps a feature vector $\mathbf{z}_{v}\in\mathbb{R}^{d+3}$ to a score vector $s_{0}(\mathbf{z}_{v}),\ldots,s_{K-1}(\mathbf{z}_{v})]$. In probabilistic models, these scores are converted into class probabilities using the softmax function:
$P(\ell_{v}=k\mid\mathbf{z}_{v};\theta)=\frac{\exp(s_{k}(\mathbf{z}_{v}))}{\sum_{j=0}^{K-1}\exp(s_{j}(\mathbf{z}_{v}))}$ (36)
while non-probabilistic models rely on decision functions or ensemble voting strategies (Hastie et al., 2009). The predicted label is selected as $\hat{\ell}_{v}=\arg\max_{k}s_{k}(\mathbf{z}_{v})$.
To train the model, we minimized the regularized loss function:
$L(\theta)=\frac{1}{N}\sum_{i=1}^{N}\ell(f(\mathbf{z}_{i};\theta),\ell_{i})+\lambda R(\theta)$ (37)
where $\ell(\cdot,\cdot)$ is a task-specific loss (e.g., cross-entropy, hinge loss), and $R(\theta)$ is a regularization term weighted by $\lambda$ to prevent overfitting (Bishop and Nasrabadi, 2006).
@@
To enhance generalizability across operating conditions, we adopted a transfer learning strategy. Let the source domain be $\mathcal{D}_{S}=\{(\mathbf{z}_{i}^{S},\ell_{i}^{S})\}_{i=1}^{N_{S}}$ and the target domain be $\mathcal{D}_{T}=\{(\mathbf{z}_{i}^{T},\ell_{i}^{T})\}_{i=1}^{N_{T}}$, where both domains share the same feature space ($\mathbb{R}^{d+3}$) and label space ($\{0,1,\ldots,K-1\}$), but differ in data distributions owing to domain shifts (e.g., FFT frequency variations from 50–100 Hz to 55–95 Hz) (Zhang et al., 2017). The model is first trained on $\mathcal{D}_{S}$ by minimizing
$L(\theta)=\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\ell(f(\mathbf{z}_{i}^{S};\theta),\ell_{i}^{S})+\lambda R(\theta)$ (38)
and then adapted to $\mathcal{D}_{T}$ by initializing $\theta_{T}^{(0)}=\theta_{S}$ and fine-tuning via:
$L(\theta)=\frac{1}{N_{T}}\sum_{i=1}^{N_{T}}\ell(f(\mathbf{z}_{i}^{T};\theta),\ell_{i}^{T})+\lambda R(\theta)$ (39)
To ensure transparency and interpretability in fault diagnosis, our framework incorporates interpretable machine learning models, including Linear Regression, Support Vector Machines (SVM), and Random Forests (Hastie et al., 2009). These models deliver robust classification performance while providing insights into decision boundaries and feature importance, making them well-suited for practical deployment in industrial settings. To comprehensively assess the predictive power of the extracted features, we will evaluate three classifiers: Linear Regression, SVM, and Random Forest.
@@ 2.6 Algorithmic Overview and Complexity Analysis
The proposed framework for multiclass fault detection processes multichannel vibration signals through five stages: feature extraction, clustering, k-nearest neighbor (KNN) graph construction, graph-theoretic feature extraction, and fault classification. The whole pipeline is defined by Algorithm 1. This framework and pipeline integrate signal processing with graph-based techniques to enable robust fault detection in complex systems.
1
Data: Raw multichannel vibration signal $S\in\mathbb{R}^{N\times\text{ch}}$, labels $\ell$ (if supervised)
Result: Predicted fault labels $\hat{\ell}$
2
3 Entropy‑Driven Segmentation
4   Estimate candidate window sizes $\Omega$ and overlap ratios $\rho$;
5 foreach $(\omega,\rho)\in\Omega\times\rho$ do
6    $\sigma\leftarrow\lfloor\omega\,(1-\rho)\rfloor$;
7      Segment $S$ into $\{X_{i}\}$ with window $\omega$ and step $\sigma$;
8      Compute entropy scores $H_{t},H_{f}$ for each segment;
9      Compute $J(\omega,\sigma)=\bigl{[}\alpha\,H_{t}+(1-\alpha)\,H_{f}\bigr{]}/\log(1+\omega)$;
10
11 $(\omega^{*},\sigma^{*})\leftarrow\arg\max J$;
12
13
14 Feature Extraction
15   Segment $S$ into $\{X_{i}\}$ using $(\omega^{*},\sigma^{*})$;
16 foreach segment $X_{i}$ do
17      Compute time‑domain stats: mean, std, skewness, kurtosis, RMS, peak;
18      Compute TKEO energy;
19      Compute FFT‑based PSD and envelope spectrum via Hilbert transform;
20      Concatenate into feature vector $F_{i}\in\mathbb{R}^{d}$;
21
22
23
24 Clustering & k‑NN Graph Construction
25   Run MiniBatch K‑Means on $\{F_{i}\}$ to obtain $K_{c}$ clusters;
26 foreach cluster $C_{j}$ do
27    $V_{j}\leftarrow\{\,i:F_{i}\in C_{j}\}$;
28      Build k‑NN graph $G_{j}=(V_{j},E_{j})$:;
29    foreach node $u\in V_{j}$ do
30         Find $k$ nearest neighbors in feature space;
31         Add weighted edges $(u,v)$ if $\mathrm{dist}(u,v)<\tau$;
32
33    if $|V_{j}|>N_{\max}$ then
34        split $C_{j}$ further;
35
36
37
38 Graph‑Theoretic Feature Computation
39   Initialize lists $L,Q,\Delta$;
40 foreach subgraph $G_{j}$ do
41      Compute average shortest path length $L_{j}$;
42      Compute modularity $Q_{j}$ w.r.t. known labels $\ell$;
43      Compute spectral gap $\Delta_{j}=\lambda_{2}(\text{Laplacian}(G_{j}))$;
44      Append $L_{j}\to L$, $Q_{j}\to Q$, $\Delta_{j}\to\Delta$;
45
46
47
48 Classification / Transfer Learning
49   Form feature matrix $Z\in\mathbb{R}^{N_{\text{total}}\times(d+3)}$ by augmenting each $F_{i}$ with $(\overline{L},\overline{Q},\overline{\Delta})$;
50 if transfer learning then
51      Pretrain on source domain $(Z_{S},\ell_{S})$ using Random Forest;
52      Fine‑tune on target domain $(Z_{T},\ell_{T})$;
53
54 else
55      Train Random Forest classifier on $(Z,\ell)$;
56
57 $\hat{\ell}\leftarrow\text{RF.predict}(Z)$;
58 return $\hat{\ell}$;
59
60
Algorithm 1 Graph‑Based Fault Diagnosis Pipeline
@@ 55      Train Random Forest classifier on $(Z,\ell)$;
Feature extraction segments a $ch$-channel vibration signal $\mathbf{S}\in\mathbb{R}^{N\times ch}$ into $n=\lfloor(N-\omega^{*})/\sigma^{*}\rfloor$ segments by using the optimal window size $\omega^{*}$ and step size $\sigma^{*}$. Each segment yields a feature vector of dimension $d=ch\cdot d_{c}$, where $d_{c}$ is the number of features per channel (e.g., $d=40$ for $ch=2$, $d_{c}=20$). The FFT, the dominant operation, has complexity $O(\omega^{*}\cdot\log\omega^{*})$ per segment per channel, resulting in a total complexity of $O(ch\cdot n\cdot\omega^{*}\cdot\log\omega^{*})$.
The $n$ feature vectors are clustered into $K_{c}$ groups using MiniBatch K-means, with complexity $O(n\cdot ch\cdot d_{c}\cdot K_{c})$, enabling efficient pattern identification. For each cluster, a KNN graph connects each node to its $k_{\text{NN}}$ nearest neighbors within a threshold $\tau$, with large clusters decomposed into subgraphs of size at most $N_{\text{max}}$. With $n/K_{c}$ nodes per cluster, the complexity is $O\left(\frac{n^{2}}{K_{c}}\left(ch\cdot d_{c}+\log(n/K_{c})\right)\right)$, dominating the framework’s cost.
Graph-theoretic features are extracted from $m_{G}=\sum_{i=1}^{K_{c}}\lceil(n/K_{c})/N_{\text{max}}\rceil$ subgraphs, each containing approximately $|V_{l}|\approx n/m_{G}$ nodes. Dijkstra’s algorithm, applied to compute all pairs of shortest paths, has a complexity of $O\left(\frac{n^{2}}{m_{G}}\log(n/m_{G})\right)$. The final classification stage, using classifiers such as Support Vector Machines or Random Forests, scales linearly with $n$, contributing lower-order terms. For a small number of channels ($ch$) and moderate $n$, the KNN graph construction with complexity $O(n^{2}/K_{c})$ often dominates. However, for large $ch$ and $n$, feature extraction with complexity $O(ch\cdot n\cdot\omega^{*}\cdot\log\omega^{*})$ becomes dominant, owing to its high computational cost.